# CLAPS: Posterior-Aware Conformal Intervals via Last-Layer Laplace

### v2

@@ 5 Experiments
We evaluate CLAPS on four standard tabular regression benchmarks spanning small to large scales and varying degrees of heterogeneity: UCI Airfoil (1,503 examples; 5 continuous input features; target is sound pressure level), OpenML kin8nm (8,192 examples; 8 continuous features; robotic arm kinematics response), UCI CASP, the Physicochemical Properties of Protein Tertiary Structure dataset (45,730 examples; 9 physicochemical features; target is RMSD), and UCI YearPredictionMSD (a fixed split with 463,715 train and 51,630 test instances; 90 audio features; target is release year). All features are numeric; we standardize inputs to zero mean and unit variance using statistics from the training split only, and we center the target during model fitting but report all metrics in the original target scale. These datasets cover regimes where last-layer epistemic variance is nontrivial in small/medium sets (Airfoil, kin8nm, CASP) and globally contracted in large-scale settings (YearPredictionMSD), allowing us to probe when posterior-aware scoring is most beneficial.
@@
All methods share the same backbone MLP to isolate the effect of uncertainty handling. For Airfoil and kin8nm, we use a 2-layer MLP with 128 hidden units per layer; for CASP and YearPredictionMSD, we use a 3-layer MLP with 256 hidden units per layer. The output head for the base predictor is a single linear neuron (scalar regression). Our method (CLAPS) applies a last-layer Laplace approximation on this linear head only, leaving the backbone fixed; competing baselines reuse the identical backbone and differ only in the head (e.g., a single point head for Baseline-CP, a positive scale head for Normalized-CP, and two quantile heads for CQR).
@@
  • Baseline-CP (CP) (lei2018distribution): A single-point head trained for squared error; calibrate absolute residuals $|y-\hat{\mu}(x)|$ on the calibration split and output $\hat{\mu}(x)\pm q_{1-\alpha}$.
  • Normalized-CP (Norm-CP) (papadopoulos2008normalized): Add a positive scale head $h(x)>0$ and use the normalized residual $|y-\hat{\mu}(x)|/h(x)$ as the score to obtain input-adaptive (heteroscedastic) widths.
  • Conformalized Quantile Regression (CQR) (romano2019conformalized): Replace the point head with two quantile heads $(\hat{q}_{\mathrm{lo}}(x),\hat{q}_{\mathrm{hi}}(x))$ and calibrate the nonnegative score $\max\{\hat{q}_{\mathrm{lo}}(x)-y,\ y-\hat{q}_{\mathrm{hi}}(x),\ 0\}$ to produce the narrowest interval consistent with split ranks.
  • Conservative Targeted Intervals (CTI) (luo2025conformal): Pretrain multiple candidate target coverages $\{1-\alpha_{k}\}$ and select at calibration time via an integer-valued rank score; typically trades width for higher reliability.
@@ Results.
Table 1: Ablation on Airfoil at $\alpha=0.10$: effect of $\lambda$ and $\hat{\sigma}^{2}$ estimator on coverage, width, and resulting $\tau$. 95% CIs are reported in the Appendix.
@@ Discussion.
CLAPS aligns the conformal nonconformity with the last-layer Laplace posterior, which in our experiments was most beneficial on small to medium tabular datasets where epistemic variance is nontrivial. In Airfoil and kin8nm, this alignment yielded noticeably narrower intervals at the same nominal coverage, consistent with the posterior-aware efficiency arguments; on the large-scale YearPrediction, last-layer epistemic variance collapsed ($\widehat{v}(x)\to\sigma^{2}$) and methods that learn input-dependent scales (Normalized-CP, CQR) achieved tighter widths. The ablation study highlights additional strengths of CLAPS: performance is remarkably insensitive to the regularization level $\lambda$ and to the choice of aleatoric estimator (both $\widehat{\sigma}^{2}_{\mathrm{EB}}$ and $\widehat{\sigma}^{2}_{\mathrm{res}}$ lead to essentially identical coverages and very similar widths), while the conformal quantile $\tau$ adjusts smoothly as $\lambda$ varies. This robustness reduces tuning burden, provides predictable behavior across datasets, and ensures graceful degradation when epistemic variance is small, while conformal calibration still guarantees exact finite-sample validity. Conceptually, scale- or quantile-based conformalizers primarily target aleatoric variability, whereas CLAPS leverages epistemic tail information captured by the last-layer posterior; this regime split clarifies when each family is preferable. Practically, CLAPS can serve as a complementary option selected by simple diagnostics such as the epistemic share $r(x)$, the posterior trace $\operatorname{tr}(\Sigma)$, or the rank correlation between $|y-\hat{\mu}(x)|$ and $\sqrt{\hat{v}(x)}$; a hybrid formulation that interpolates between a posterior-aware term and a learned scale term using $r$ is a promising direction for future work.
@@ Appendix C Extended Ablations and Sensitivity Analyses
This section provides the full ablation grids that underlie the main text. Table 6 reports, for each dataset (Airfoil, kin8nm, CASP, YearPredictionMSD), the core quantities across regularization levels $\lambda\!\in\!\{0.1,0.3,1,3,10\}$ and two aleatoric estimators $\widehat{\sigma}^{2}_{\mathrm{EB}}$ and $\widehat{\sigma}^{2}_{\mathrm{res}}$: empirical coverage (Cov), average interval width (Wid), the selected conformal quantile $\tau$, and the fitted noise level $\widehat{\sigma}^{2}$. Table 7 complements these results with 95% confidence intervals for coverage and width only, enabling a quantitative assessment of the small empirical differences observed among settings.
@@
Several consistent patterns emerge. First, coverage remains essentially invariant across $\lambda$ and across the two choices of $\widehat{\sigma}^{2}$, confirming that the conformal rank rule renders CLAPS insensitive to reasonable regularization and noise-estimation choices. Second, widths change only modestly with $\lambda$, while $\tau$ adapts smoothly so that nominal coverage is preserved; this monotone adjustment is most visible on Airfoil and CASP. Third, $\widehat{\sigma}^{2}_{\mathrm{EB}}$ and $\widehat{\sigma}^{2}_{\mathrm{res}}$ lead to nearly identical operating points across all datasets, suggesting that either can be used as a practical default without sacrificing efficiency. Finally, on the large-scale YearPredictionMSD, the reported dimensionality $D$ of the last-layer features and the calibration size $n$ (annotated in the table) coincide with a regime where the epistemic share is small; the grids illustrate that CLAPS degrades gracefully in such cases, maintaining coverage while producing widths in line with the learned-scale baselines reported in the main text. Together, Tables 6 and 7 document that CLAPS exhibits low tuning burden and predictable behavior across datasets and hyperparameters, while retaining exact finite-sample validity.
$\lambda$ Estimator Cover Wid $\tau$ $\widehat{\sigma}^{2}$ 0.1 $\hat{\sigma}^{2}_{\mathrm{EB}}$ 0.9007 7.7100 0.036598 0.0879627 $\hat{\sigma}^{2}_{\mathrm{res}}$ 0.9007 7.7106 0.030127 0.0799660 0.3 $\hat{\sigma}^{2}_{\mathrm{EB}}$ 0.8742 7.5557 0.035150 0.0831758 $\hat{\sigma}^{2}_{\mathrm{res}}$ 0.8742 7.5557 0.035150 0.0831758 1.0 $\hat{\sigma}^{2}_{\mathrm{EB}}$ 0.9007 7.7842 0.034182 0.0880946 $\hat{\sigma}^{2}_{\mathrm{res}}$ 0.9007 7.7842 0.034182 0.0880946 3.0 $\hat{\sigma}^{2}_{\mathrm{EB}}$ 0.9007 7.9756 0.033639 0.0932376 $\hat{\sigma}^{2}_{\mathrm{res}}$ 0.9007 7.9756 0.033639 0.0932376 10.0 $\hat{\sigma}^{2}_{\mathrm{EB}}$ 0.9073 8.2366 0.025537 0.0890167 $\hat{\sigma}^{2}_{\mathrm{res}}$ 0.9073 8.2366 0.031979 0.0989075
$\lambda$ Estimator Cov Wid $\tau$ $\widehat{\sigma}^{2}$ 0.1 $\hat{\sigma}^{2}_{\mathrm{EB}}$ 0.9037 0.2669 0.040228 0.0819326 $\hat{\sigma}^{2}_{\mathrm{res}}$ 0.9037 0.2669 0.040228 0.0819326 0.3 $\hat{\sigma}^{2}_{\mathrm{EB}}$ 0.9049 0.2663 0.040519 0.0819418 $\hat{\sigma}^{2}_{\mathrm{res}}$ 0.9049 0.2663 0.040519 0.0819418 1.0 $\hat{\sigma}^{2}_{\mathrm{EB}}$ 0.9024 0.2670 0.040146 0.0819847 $\hat{\sigma}^{2}_{\mathrm{res}}$ 0.9024 0.2670 0.040146 0.0819847 3.0 $\hat{\sigma}^{2}_{\mathrm{EB}}$ 0.9049 0.2702 0.038544 0.0821850 $\hat{\sigma}^{2}_{\mathrm{res}}$ 0.9049 0.2702 0.038544 0.0821850 10.0 $\hat{\sigma}^{2}_{\mathrm{EB}}$ 0.9061 0.2700 0.039473 0.0831243 $\hat{\sigma}^{2}_{\mathrm{res}}$ 0.9061 0.2700 0.039473 0.0831243
$\lambda$ Estimator Cov Wid $\tau$ $\widehat{\sigma}^{2}$ 0.1 $\hat{\sigma}^{2}_{\mathrm{EB}}$ 0.9033 13.8806 0.035899 0.394733 $\hat{\sigma}^{2}_{\mathrm{res}}$ 0.9033 13.8806 0.035899 0.394733 0.3 $\hat{\sigma}^{2}_{\mathrm{EB}}$ 0.9031 13.8899 0.035849 0.395051 $\hat{\sigma}^{2}_{\mathrm{res}}$ 0.9031 13.8899 0.035849 0.395051 1.0 $\hat{\sigma}^{2}_{\mathrm{EB}}$ 0.9033 13.9118 0.035769 0.396034 $\hat{\sigma}^{2}_{\mathrm{res}}$ 0.9033 13.9118 0.035769 0.396034 3.0 $\hat{\sigma}^{2}_{\mathrm{EB}}$ 0.9058 13.9439 0.035740 0.398011 $\hat{\sigma}^{2}_{\mathrm{res}}$ 0.9058 13.9439 0.035740 0.398011 10.0 $\hat{\sigma}^{2}_{\mathrm{EB}}$ 0.9066 13.9214 0.036450 0.401143 $\hat{\sigma}^{2}_{\mathrm{res}}$ 0.9066 13.9214 0.036450 0.401143
$\lambda$ Estimator Cov Wid $\tau$ $\widehat{\sigma}^{2}$ 0.1 $\hat{\sigma}^{2}_{\mathrm{EB}}$ 0.9019 27.5707 0.043697 0.545348 $\hat{\sigma}^{2}_{\mathrm{res}}$ 0.9019 27.5707 0.043697 0.545348 0.3 $\hat{\sigma}^{2}_{\mathrm{EB}}$ 0.9017 27.5692 0.043708 0.545369 $\hat{\sigma}^{2}_{\mathrm{res}}$ 0.9017 27.5692 0.043708 0.545369 1.0 $\hat{\sigma}^{2}_{\mathrm{EB}}$ 0.9020 27.5889 0.043604 0.545434 $\hat{\sigma}^{2}_{\mathrm{res}}$ 0.9020 27.5889 0.043604 0.545434 3.0 $\hat{\sigma}^{2}_{\mathrm{EB}}$ 0.9017 27.5727 0.043713 0.545558 $\hat{\sigma}^{2}_{\mathrm{res}}$ 0.9017 27.5727 0.043713 0.545558 10.0 $\hat{\sigma}^{2}_{\mathrm{EB}}$ 0.9016 27.6128 0.043516 0.545810 $\hat{\sigma}^{2}_{\mathrm{res}}$ 0.9016 27.6128 0.043516 0.545810
@@
Table 6: Ablation summary at $\alpha=0.10$ across four datasets. Columns report finite-sample coverage, mean interval width, rank-derived threshold $\tau$, and the fitted noise variance $\widehat{\sigma}^{2}$.
@@ D.1 Preliminaries and Notation
$w\mid\mathcal{D}_{\mathrm{tr}}\sim\mathcal{N}\!\big(w_{\mathrm{MAP}},\,\Sigma\big),\qquad\Sigma\;=\;\big(\lambda I_{d}+\sigma^{-2}\Phi^{\top}\Phi\big)^{-1},\quad\Phi=\big[\phi(x_{i})^{\top}\big]_{i=1}^{n_{\mathrm{tr}}}.$
@@
Numerical identity (Cholesky form). If $LL^{\top}=\lambda I_{d}+\sigma^{-2}\Phi^{\top}\Phi$ with $L$ lower triangular, then
@@
Assumption 1. Conditional on $\mathcal{D}_{\mathrm{tr}}$, the multiset $\mathcal{D}_{\mathrm{cal}}$ and the test pair $(x_{\ast},y_{\ast})$ are exchangeable draws from the same data-generating process. This assumption is only invoked for the marginal validity of split-conformal calibration.
@@
We write $\mathbb{P}[\cdot]$ and $\mathbb{E}[\cdot]$ for probability and expectation over the randomness of $(\mathcal{D}_{\mathrm{cal}},x_{\ast},y_{\ast})$ under Assumption 1. The Lebesgue measure (length) of an interval $A\subset\mathbb{R}$ is denoted by $|A|$. All asymptotics ($n_{\mathrm{tr}}\!\to\!\infty$) keep the backbone architecture fixed unless explicitly stated.
@@
Consider the augmented (m+1)(m\!+\!1) scores consisting of the $m$ calibration scores and the independent test score $s_{\ast}:=s(x_{\ast},y_{\ast})$. By Assumption 1 (exchangeability of $(x_{\ast},y_{\ast})$ with the calibration pairs conditional on $\mathcal{D}_{\mathrm{tr}}$), the multiset $\{s_{1},\dots,s_{m},s_{\ast}\}$ is exchangeable. Let $R$ denote the (randomized, measurable) rank of $s_{\ast}$ among these (m+1)(m\!+\!1) scores in ascending order (ties are broken by an independent uniform randomization or any fixed measurable rule). Exchangeability implies
@@
Write $\Sigma_{n}=(\lambda I_{d}+\sigma^{-2}\Phi_{n}^{\top}\Phi_{n})^{-1}$ for the posterior covariance built from the $n=n_{\mathrm{tr}}$ training points, where $\Phi_{n}=[\phi(x_{i})^{\top}]_{i=1}^{n}\in\mathbb{R}^{n\times d}$ and $d$ is the (fixed) backbone width. Assume throughout that $\mathbb{E}[\phi(x)\phi(x)^{\top}]=:\Gamma$ exists with $\lambda_{\min}(\Gamma)>0$ and that $\|\phi(x)\|_{2}$ has finite second moment (e.g., sub-Gaussian or bounded features suffice).
Step 1: Spectral growth of $\Phi_{n}^{\top}\Phi_{n}$.
@@
$\frac{1}{n}\,\Phi_{n}^{\top}\Phi_{n}\;\xrightarrow[n\to\infty]{\text{a.s.}}\;\Gamma.$
@@
$\lambda_{\min}(\Phi_{n}^{\top}\Phi_{n})\;=\;n\,\lambda_{\min}\!\left(\tfrac{1}{n}\Phi_{n}^{\top}\Phi_{n}\right)\;\geq\;n\,\frac{\lambda_{\min}(\Gamma)}{2}\qquad\text{for all sufficiently large }n.$
@@
$\lambda_{\min}\!\big(\lambda I_{d}+\sigma^{-2}\Phi_{n}^{\top}\Phi_{n}\big)\;\geq\;\lambda+\frac{\lambda_{\min}(\Gamma)}{2\sigma^{2}}\,n\;\xrightarrow{n\to\infty}\;\infty,$
@@
$\|\Sigma_{n}\|_{\mathrm{op}}=\big\|(\lambda I_{d}+\sigma^{-2}\Phi_{n}^{\top}\Phi_{n})^{-1}\big\|_{\mathrm{op}}=\frac{1}{\lambda_{\min}(\lambda I_{d}+\sigma^{-2}\Phi_{n}^{\top}\Phi_{n})}\;\xrightarrow{n\to\infty}\;0.$
@@
The conclusion holds for fixed backbone width $d$ and nondegenerate feature covariance $\Gamma\succ 0$. If $d$ grows with $n$ in such a way that $\lambda_{\min}(\tfrac{1}{n}\Phi_{n}^{\top}\Phi_{n})$ fails to be bounded away from 0, or if features are degenerate ($\Gamma\succeq 0$ with zero eigenvalues along directions used by $\phi(x)$), then $\|\Sigma_{n}\|_{\mathrm{op}}\to 0$ need not hold uniformly, and residual epistemic variance may persist. ∎

### v3

@@ 5 Experiments
We evaluate CLAPS on four standard tabular regression benchmarks spanning small to large scales and varying degrees of heterogeneity: UCI Airfoil (1,503 examples; 5 continuous input features; target is sound pressure level), OpenML kin8nm (8,192 examples; 8 continuous features; robotic arm kinematics response), UCI CASP, the Physicochemical Properties of Protein Tertiary Structure dataset (45,730 examples; 9 physicochemical features; target is RMSD), and UCI YearPredictionMSD (a fixed split with 463,715 train and 51,630 test instances; 90 audio features; target is release year). All features are numeric; we standardize input features and the target using statistics computed on the training split only (zero mean and unit variance), and report all metrics on the original target scale via the inverse transform. These datasets cover regimes where last-layer epistemic variance is nontrivial in small/medium sets (Airfoil, kin8nm, CASP) and globally contracted in large-scale settings (YearPredictionMSD), allowing us to probe when posterior-aware scoring is most beneficial.
@@
All methods share the same backbone MLP to control for representation capacity and focus the comparison on the uncertainty / interval construction layer. For Airfoil and kin8nm, we use a 2-layer MLP with 128 hidden units per layer; for CASP and YearPredictionMSD, we use a 3-layer MLP with 256 hidden units per layer. The base predictor uses a single linear neuron (scalar regression). Our method (CLAPS) applies a last-layer Laplace approximation only to this linear head, leaving the backbone unchanged. Baselines reuse the identical backbone but may use different output heads required by each method (e.g., a single point head for Baseline-CP, an additional positive scale head for Normalized-CP, and two quantile heads for CQR). Accordingly, point-prediction metrics (e.g., MAE) can be influenced by head-specific training objectives; we therefore emphasize coverage and interval efficiency as the primary comparison criteria.
@@
  • Baseline-CP (CP) (Lei et al., 2018): A single-point head trained for squared error; calibrate absolute residuals $|y-\hat{\mu}(x)|$ on the calibration split and output $\hat{\mu}(x)\pm q_{1-\alpha}$.
  • Normalized-CP (Norm-CP) (Papadopoulos et al., 2008): Add a positive scale head $h(x)>0$ and use the normalized residual $|y-\hat{\mu}(x)|/h(x)$ as the score to obtain input-adaptive (heteroscedastic) widths.
  • Conformalized Quantile Regression (CQR) (Romano et al., 2019): Replace the point head with two quantile heads $(\hat{q}_{\mathrm{lo}}(x),\hat{q}_{\mathrm{hi}}(x))$ and calibrate the nonnegative score $\max\{\hat{q}_{\mathrm{lo}}(x)-y,\ y-\hat{q}_{\mathrm{hi}}(x),\ 0\}$ to produce the narrowest interval consistent with split ranks.
  • Conservative Targeted Intervals (CTI) (Luo and Zhou, 2025): Pretrain multiple candidate target coverages $\{1-\alpha_{k}\}$ and select at calibration time via an integer-valued rank score; typically trades width for higher reliability.
@@ Results.
Table 1: Ablation on Airfoil at $\alpha=0.10$: effect of $\lambda$ and $\hat{\sigma}^{2}$ estimator on coverage, width, and resulting $t$. 95% CIs are reported in the Appendix.
@@ Discussion.
CLAPS aligns the conformal nonconformity with the last-layer Laplace posterior, which in our experiments was most beneficial on small to medium tabular datasets where epistemic variance is nontrivial. In Airfoil and kin8nm, this alignment yielded noticeably narrower intervals at the same nominal coverage, consistent with the posterior-aware efficiency arguments; on the large-scale YearPrediction, last-layer epistemic variance collapsed ($\widehat{v}(x)\to\sigma^{2}$) and methods that learn input-dependent scales (Normalized-CP, CQR) achieved tighter widths. The ablation study highlights additional strengths of CLAPS: performance is remarkably insensitive to the regularization level $\lambda$ and to the choice of aleatoric estimator (both $\widehat{\sigma}^{2}_{\mathrm{EB}}$ and $\widehat{\sigma}^{2}_{\mathrm{res}}$ lead to essentially identical coverages and very similar widths), while the conformal quantile $t$ adjusts smoothly as $\lambda$ varies. This robustness reduces tuning burden, provides predictable behavior across datasets, and ensures graceful degradation when epistemic variance is small, while conformal calibration still guarantees exact finite-sample validity. Conceptually, scale- or quantile-based conformalizers primarily target aleatoric variability, whereas CLAPS leverages epistemic tail information captured by the last-layer posterior; this regime split clarifies when each family is preferable. Practically, CLAPS can serve as a complementary option selected by simple diagnostics such as the epistemic share $r(x)$, the posterior trace $\operatorname{tr}(\Sigma)$, or the rank correlation between $|y-\hat{\mu}(x)|$ and $\sqrt{\hat{v}(x)}$; a hybrid formulation that interpolates between a posterior-aware term and a learned scale term using $r$ is a promising direction for future work.
@@ Appendix C Extended Ablations and Sensitivity Analyses
This section provides the full ablation grids that underlie the main text. Table 6 reports, for each dataset (Airfoil, kin8nm, CASP, YearPredictionMSD), the core quantities across regularization levels $\lambda\!\in\!\{0.1,0.3,1,3,10\}$ and two aleatoric estimators $\widehat{\sigma}^{2}_{\mathrm{EB}}$ and $\widehat{\sigma}^{2}_{\mathrm{res}}$: empirical coverage (Cov), average interval width (Wid), the selected conformal quantile $t$, and the fitted noise level $\widehat{\sigma}^{2}$. Table 7 complements these results with 95% confidence intervals for coverage and width only, enabling a quantitative assessment of the small empirical differences observed among settings.
@@
Several consistent patterns emerge. First, coverage remains essentially invariant across $\lambda$ and across the two choices of $\widehat{\sigma}^{2}$, confirming that the conformal rank rule renders CLAPS insensitive to reasonable regularization and noise-estimation choices. Second, widths change only modestly with $\lambda$, while $t$ adapts smoothly so that nominal coverage is preserved; this monotone adjustment is most visible on Airfoil and CASP. Third, $\widehat{\sigma}^{2}_{\mathrm{EB}}$ and $\widehat{\sigma}^{2}_{\mathrm{res}}$ lead to nearly identical operating points across all datasets, suggesting that either can be used as a practical default without sacrificing efficiency. Finally, on the large-scale YearPredictionMSD, the reported dimensionality $D$ of the last-layer features and the calibration size $n$ (annotated in the table) coincide with a regime where the epistemic share is small; the grids illustrate that CLAPS degrades gracefully in such cases, maintaining coverage while producing widths in line with the learned-scale baselines reported in the main text. Together, Tables 6 and 7 document that CLAPS exhibits low tuning burden and predictable behavior across datasets and hyperparameters, while retaining exact finite-sample validity.
$\lambda$ Estimator Cover Wid $t$ $\widehat{\sigma}^{2}$ 0.1 $\hat{\sigma}^{2}_{\mathrm{EB}}$ 0.9007 7.7100 0.036598 0.0879627 $\hat{\sigma}^{2}_{\mathrm{res}}$ 0.9007 7.7106 0.030127 0.0799660 0.3 $\hat{\sigma}^{2}_{\mathrm{EB}}$ 0.8742 7.5557 0.035150 0.0831758 $\hat{\sigma}^{2}_{\mathrm{res}}$ 0.8742 7.5557 0.035150 0.0831758 1.0 $\hat{\sigma}^{2}_{\mathrm{EB}}$ 0.9007 7.7842 0.034182 0.0880946 $\hat{\sigma}^{2}_{\mathrm{res}}$ 0.9007 7.7842 0.034182 0.0880946 3.0 $\hat{\sigma}^{2}_{\mathrm{EB}}$ 0.9007 7.9756 0.033639 0.0932376 $\hat{\sigma}^{2}_{\mathrm{res}}$ 0.9007 7.9756 0.033639 0.0932376 10.0 $\hat{\sigma}^{2}_{\mathrm{EB}}$ 0.9073 8.2366 0.025537 0.0890167 $\hat{\sigma}^{2}_{\mathrm{res}}$ 0.9073 8.2366 0.031979 0.0989075
$\lambda$ Estimator Cov Wid $t$ $\widehat{\sigma}^{2}$ 0.1 $\hat{\sigma}^{2}_{\mathrm{EB}}$ 0.9037 0.2669 0.040228 0.0819326 $\hat{\sigma}^{2}_{\mathrm{res}}$ 0.9037 0.2669 0.040228 0.0819326 0.3 $\hat{\sigma}^{2}_{\mathrm{EB}}$ 0.9049 0.2663 0.040519 0.0819418 $\hat{\sigma}^{2}_{\mathrm{res}}$ 0.9049 0.2663 0.040519 0.0819418 1.0 $\hat{\sigma}^{2}_{\mathrm{EB}}$ 0.9024 0.2670 0.040146 0.0819847 $\hat{\sigma}^{2}_{\mathrm{res}}$ 0.9024 0.2670 0.040146 0.0819847 3.0 $\hat{\sigma}^{2}_{\mathrm{EB}}$ 0.9049 0.2702 0.038544 0.0821850 $\hat{\sigma}^{2}_{\mathrm{res}}$ 0.9049 0.2702 0.038544 0.0821850 10.0 $\hat{\sigma}^{2}_{\mathrm{EB}}$ 0.9061 0.2700 0.039473 0.0831243 $\hat{\sigma}^{2}_{\mathrm{res}}$ 0.9061 0.2700 0.039473 0.0831243
$\lambda$ Estimator Cov Wid $t$ $\widehat{\sigma}^{2}$ 0.1 $\hat{\sigma}^{2}_{\mathrm{EB}}$ 0.9033 13.8806 0.035899 0.394733 $\hat{\sigma}^{2}_{\mathrm{res}}$ 0.9033 13.8806 0.035899 0.394733 0.3 $\hat{\sigma}^{2}_{\mathrm{EB}}$ 0.9031 13.8899 0.035849 0.395051 $\hat{\sigma}^{2}_{\mathrm{res}}$ 0.9031 13.8899 0.035849 0.395051 1.0 $\hat{\sigma}^{2}_{\mathrm{EB}}$ 0.9033 13.9118 0.035769 0.396034 $\hat{\sigma}^{2}_{\mathrm{res}}$ 0.9033 13.9118 0.035769 0.396034 3.0 $\hat{\sigma}^{2}_{\mathrm{EB}}$ 0.9058 13.9439 0.035740 0.398011 $\hat{\sigma}^{2}_{\mathrm{res}}$ 0.9058 13.9439 0.035740 0.398011 10.0 $\hat{\sigma}^{2}_{\mathrm{EB}}$ 0.9066 13.9214 0.036450 0.401143 $\hat{\sigma}^{2}_{\mathrm{res}}$ 0.9066 13.9214 0.036450 0.401143
$\lambda$ Estimator Cov Wid $t$ $\widehat{\sigma}^{2}$ 0.1 $\hat{\sigma}^{2}_{\mathrm{EB}}$ 0.9019 27.5707 0.043697 0.545348 $\hat{\sigma}^{2}_{\mathrm{res}}$ 0.9019 27.5707 0.043697 0.545348 0.3 $\hat{\sigma}^{2}_{\mathrm{EB}}$ 0.9017 27.5692 0.043708 0.545369 $\hat{\sigma}^{2}_{\mathrm{res}}$ 0.9017 27.5692 0.043708 0.545369 1.0 $\hat{\sigma}^{2}_{\mathrm{EB}}$ 0.9020 27.5889 0.043604 0.545434 $\hat{\sigma}^{2}_{\mathrm{res}}$ 0.9020 27.5889 0.043604 0.545434 3.0 $\hat{\sigma}^{2}_{\mathrm{EB}}$ 0.9017 27.5727 0.043713 0.545558 $\hat{\sigma}^{2}_{\mathrm{res}}$ 0.9017 27.5727 0.043713 0.545558 10.0 $\hat{\sigma}^{2}_{\mathrm{EB}}$ 0.9016 27.6128 0.043516 0.545810 $\hat{\sigma}^{2}_{\mathrm{res}}$ 0.9016 27.6128 0.043516 0.545810
@@
Table 6: Ablation summary at $\alpha=0.10$ across four datasets. Columns report finite-sample coverage, mean interval width, rank-derived threshold $t$, and the fitted noise variance $\widehat{\sigma}^{2}$.
@@ D.1 Preliminaries and Notation
$w\mid\mathcal{D}_{\mathrm{tr}}\sim\mathcal{N}\!\big(w_{\mathrm{MAP}},\,\Sigma\big),\qquad\Sigma\;=\;\big(\lambda I_{d}+\sigma^{-2}\mathbf{H}^{\top}\mathbf{H}\big)^{-1},\quad\mathbf{H}=\big[\phi(x_{i})^{\top}\big]_{i=1}^{n_{\mathrm{tr}}}.$
@@
Numerical identity (Cholesky form). If $LL^{\top}=\lambda I_{d}+\sigma^{-2}\mathbf{H}^{\top}\mathbf{H}$ with $L$ lower triangular, then
@@
Assumption 4.3. Conditional on $\mathcal{D}_{\mathrm{tr}}$, the multiset $\mathcal{D}_{\mathrm{cal}}$ and the test pair $(x_{\ast},y_{\ast})$ are exchangeable draws from the same data-generating process. This assumption is only invoked for the marginal validity of split-conformal calibration.
@@
We write $\mathbb{P}[\cdot]$ and $\mathbb{E}[\cdot]$ for probability and expectation over the randomness of $(\mathcal{D}_{\mathrm{cal}},x_{\ast},y_{\ast})$ under Assumption 4.3. The Lebesgue measure (length) of an interval $A\subset\mathbb{R}$ is denoted by $|A|$. All asymptotics ($n_{\mathrm{tr}}\!\to\!\infty$) keep the backbone architecture fixed unless explicitly stated.
@@
Consider the augmented (m+1)(m\!+\!1) scores consisting of the $m$ calibration scores and the independent test score $s_{\ast}:=s(x_{\ast},y_{\ast})$. By Assumption 4.3 (exchangeability of $(x_{\ast},y_{\ast})$ with the calibration pairs conditional on $\mathcal{D}_{\mathrm{tr}}$), the multiset $\{s_{1},\dots,s_{m},s_{\ast}\}$ is exchangeable. Let $R$ denote the (randomized, measurable) rank of $s_{\ast}$ among these (m+1)(m\!+\!1) scores in ascending order (ties are broken by an independent uniform randomization or any fixed measurable rule). Exchangeability implies
@@
Write $\Sigma_{n}=(\lambda I_{d}+\sigma^{-2}\mathbf{H}_{n}^{\top}\mathbf{H}_{n})^{-1}$ for the posterior covariance built from the $n=n_{\mathrm{tr}}$ training points, where $\mathbf{H}_{n}=[\phi(x_{i})^{\top}]_{i=1}^{n}\in\mathbb{R}^{n\times d}$ and $d$ is the (fixed) backbone width. Assume throughout that $\mathbb{E}[\phi(x)\phi(x)^{\top}]=:\Gamma$ exists with $\lambda_{\min}(\Gamma)>0$ and that $\|\phi(x)\|_{2}$ has finite second moment (e.g., sub-Gaussian or bounded features suffice).
Step 1: Spectral growth of $\mathbf{H}_{n}^{\top}\mathbf{H}_{n}$.
@@
$\frac{1}{n}\,\mathbf{H}_{n}^{\top}\mathbf{H}_{n}\;\xrightarrow[n\to\infty]{\text{a.s.}}\;\Gamma.$
@@
$\lambda_{\min}(\mathbf{H}_{n}^{\top}\mathbf{H}_{n})\;=\;n\,\lambda_{\min}\!\left(\tfrac{1}{n}\mathbf{H}_{n}^{\top}\mathbf{H}_{n}\right)\;\geq\;n\,\frac{\lambda_{\min}(\Gamma)}{2}\qquad\text{for all sufficiently large }n.$
@@
$\lambda_{\min}\!\big(\lambda I_{d}+\sigma^{-2}\mathbf{H}_{n}^{\top}\mathbf{H}_{n}\big)\;\geq\;\lambda+\frac{\lambda_{\min}(\Gamma)}{2\sigma^{2}}\,n\;\xrightarrow{n\to\infty}\;\infty,$
@@
$\|\Sigma_{n}\|_{\mathrm{op}}=\big\|(\lambda I_{d}+\sigma^{-2}\mathbf{H}_{n}^{\top}\mathbf{H}_{n})^{-1}\big\|_{\mathrm{op}}=\frac{1}{\lambda_{\min}(\lambda I_{d}+\sigma^{-2}\mathbf{H}_{n}^{\top}\mathbf{H}_{n})}\;\xrightarrow{n\to\infty}\;0.$
@@
The conclusion holds for fixed backbone width $d$ and nondegenerate feature covariance $\Gamma\succ 0$. If $d$ grows with $n$ in such a way that $\lambda_{\min}(\tfrac{1}{n}\mathbf{H}_{n}^{\top}\mathbf{H}_{n})$ fails to be bounded away from 0, or if features are degenerate ($\Gamma\succeq 0$ with zero eigenvalues along directions used by $\phi(x)$), then $\|\Sigma_{n}\|_{\mathrm{op}}\to 0$ need not hold uniformly, and residual epistemic variance may persist. ∎